Loren drove 200 miles at a certain rate, and his wife, Lois, drove 100 miles at a rate 10 mph slower. If Loren had driven for the entire trip, they would have arrived 30 minutes sooner. What was Loren's rate?
Answer by richwmiller(17219)

r*t=d
n=loren
s=lois
t= loren's time
T=lois's time
n*t=200
s*T=100
s=n-10
300=n*(t+T-1/2)
4 equations 4 unknowns
n = 50, s = 40, t = 4, T = 5/2
loren's speed was 50 mph
check
loren
50*4=200
lois
40*5/2=100
loren whole trip
50*(4+5/2-1/2)=300
[50*6=300]
ok